How do I solve this problem for Solving Systems of Linear Equations by Addition:
-10x-y= -3
-2x+y= -9

Answer by Alan3354(69443)

How do I solve this problem for Solving Systems of Linear Equations by Addition:
-10x-y= -3
-2x +y= -9
Since the y terms are both 1, add the equations.
-10x-y= -3
-2x +y= -9
------------- Add
-12x = -12
x = 1
------
-2*1 + y = -9
y = -7